OpenStudy (anonymous):

write the equivalent equation that expresses y in terms of x. 2x+2x-y=3y and 2x(2-x)=x-y can someone please help with this..

OpenStudy (anonymous):

For the first one: You need to firstly collect the like terms. \[2x+2x-y=3y\] \[4x=3y+y\] \[4x=4y\] Then make y the subject to express y in terms of x. \[4y=4x\] \[y=4x/4\] \[\therefore y=x\] For the second one: You need to expand the brackets. \[2x(2-x)=x-y\] \[4x-2x^2=x-y\] Then get y by itself. \[4x-2x^2+y=x\] \[y=2x^2+x-4x\] Finally collect the like terms. \[\therefore y=2x^2-3x\]